OpenStudy (anonymous):

in the sequence of non-zero numbers x, x^2, x^3, x^4, ..., .if from the third term on, each term is the sum of the preceding 2 terms, what are all the possible values of x

myininaya (myininaya):

I'm guess those superscripts are definitely not exponents?

OpenStudy (anonymous):

its x raised to 1,2,3,4... sory I don't know how to sophistically type them on my pc keyboard :)

myininaya (myininaya):

so they are exponents

myininaya (myininaya):

\[x,x^2,x^3,x^4,x^5,...\]

myininaya (myininaya):

but if that is so then this is also a geometric series

myininaya (myininaya):

but anyways you are given that \[x^3+x^4=x^5\] or ... \[x^{n-2}+x^{n-1}=x^{n} \text{ for } n \ge 5\]

OpenStudy (anonymous):

i found n >or = 3 but 5 makes sense

myininaya (myininaya):

well it said for the third term and on...

myininaya (myininaya):

so it sounded like x^1+x^2 doesn't imply the third term is x^3

myininaya (myininaya):

wait maybe it does mean that

myininaya (myininaya):

maybe you are right

myininaya (myininaya):

\[x^{n-2}+x^{n-1}=x^{n} \text{ for } n \ge 3 \]

OpenStudy (anonymous):

that' s what i have so far

myininaya (myininaya):

but if we just look at the first case n=3 \[x+x^2=x^3 \] we should be able to solve this equation for x

myininaya (myininaya):

we already know x is not 0 so divide both sides by x and you have a quadratic to solve

OpenStudy (anonymous):

omg! so 1+x=x^2.... yeah that makes sense! thanks!

myininaya (myininaya):

you don't get pretty values

OpenStudy (anonymous):

quadratic formula! ^_^ hehe its fine for me

myininaya (myininaya):

have fun

OpenStudy (anonymous):

thanks!